OpenStudy (anonymous):

Evaluate 16 - 20 - (-8) - 9.

OpenStudy (texaschic101):

16 - 20 - (-8) - 9 = 16 - 20 + 8 - 9 = 24 - 29 = -5

OpenStudy (anonymous):

Evaluate -3 3.

OpenStudy (anonymous):

If f(n) = n 2 - n, then f(-4) is _____.